Worksheet 18: Case Study—Assessment of the Nutritional Status of the Elderly

John is a 78-year-old male, who weighs 146#, is 5’11” and is considered sedentary. He is a sometimes homeless veteran and is missing many teeth from years of no dental care. John receives a monthly retirement check but he has great difficulty meeting his monthly bills, let alone buying enough food, and this has caused many evictions. When he can afford them, John has his prescriptions filled for a diuretic (for high blood pressure) and a statin (for high cholesterol). John isn’t sure what his current weight is, but he has had to tighten his belt to the last notch over the past eight months. He has no family that he is still in contact with and he eats all of his meals alone. At a neighbor’s insistence, John is visiting the local senior center to eat lunch for the first time. The day he visits, each person is having a free nutrition screen completed by a program staff member.

Questions:

1.       Using the “Determine Your Nutritional Health” checklist (p. 464, 5th ed, 459, 6th ed.), assess John’s nutritional status. What is his score?

2.       Using the Mifflin-St. Jeor formula (pg. 415, 5th ed, 409, 6th ed.), calculate John’s energy needs. Show your work

3.       Determine John’s estimated protein requirements. Show your work.

4.       John has difficulty chewing due to missing many teeth, so he avoids eating many types of meats. What suggestions do you have to help him consume an adequate amount of protein?

5.       Determine John’s estimated fluid requirements.

6.     John’s usual food intake is as follows: coffee, pastry, tuna salad on white bread, cola, potato chips, canned beans, hot dogs, coffee, and pie.

What suggestions do you have to help him reduce his risk for chronic disease using dietary modifications?
Answer #1

2. Mifflin -St.Jeor formula :-

For Men : 10 x weight (kg) + 6.25 x height (cm) – 5 x Age + 5

Weight of the patient is 146lb = 146×0.45= 65.7kg

Height of the patient is 5'11" = 180.34cm

Age is 78

Formula = 10×65.7+ 6.25× 180.34 - 5× 78 +5

= 657+ 1126.8 - 390+ 5

= 657+ 736.8 +5

= 1398.8kcal is the energy need of the patient

3. John's protein requirement is 0.8( normal protein required for a Patient ) × 65.7

= 52.56 gm per day

4. Some of the measures to consume adequate amount of protein are :-

- drinking soups of mutton ,chicken

- having dairy products

- including legumes pulses ,oats etc

- eating nuts such as almond

- in vegetables eating hemp seeds ,black beans etc

5..Fluid requirement :-

For > 20 kg = 1500 mL + [weight (kg) x 20 ml/kg/day]

= 1500ml+ (65.7× 20)

= 2814 ml